OpenStudy (anonymous):

The ages of the eight-member school board are listed below. 28, 40, 34, 46, 60, 34, 51, 57 If A is equal to the median age of the group and B is equal to the mode of the group/s age, what is the value of A-B?

OpenStudy (ajprincess):

First arrange the data given in ascending order. then find the (n+1)/2th data. n-number of data available.:) it will be the median. mode is the data that occurs many times.

OpenStudy (anonymous):

mode=34 ----------(B) median=43----------(A) A-B= 43-34= 9

OpenStudy (ajprincess):

Does that help? @jenn_jenn94